# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ 4.2 Empirical Application
Method 1 Method 2 RMSFE (ratio) Selected Variables (Av.)
$h$=12 $h=24$ $h$=12 $h=24$
ARMAr-LAS LAS 0.69*** 0.82* 6.0||||67.9 6.2||||60.9
ARMAr-LAS GLS-LAS 0.66*** 0.61*** 6.0||||3.5 6.2||||3.8
[…]
ARMAr-LAS FarSel 0.71*** 0.73*** 6.0||||77.2 6.2||||72.5
ARMAr-LAS AR($p$) 0.94 0.89*
LAS AR($p$) 1.36 1.08
GLS-LAS AR($p$) 1.43 1.44
ARDL-LAS AR($p$) 1.53 1.07
FarSel AR($p$) 1.32 1.21
  • Note: For AR($p$) coefficients are estimated using the R function lm. For ARMAr-LAS estimated residuals are obtained by means of an ARMA($p_{i},q_{i}$) filter. The penalty parameter $\lambda$ is selected with BIC using the R package HDeconometrics.
Figure 3 summarizes patterns of selected predictors over time for LAS and ARMAr-LAS. The heatmaps represent the number of selected variables categorized according to the nine main domains (see above). LAS selects predictors largely, though not exclusively, from the domains ICS, M&IR and HUR. ARMAr-LAS is more targeted, selecting predictors almost exclusively in the HCPI domain. Note, however, that in a few instances (3 for $h=12$ and 2 for $h=24$) ARMAr-LAS does select many more predictors across multiple groups. Interestingly, these correspond to the period of the financial crisis (between the end of 2008 and the beginning of 2010), when negative shocks in some of the variables under consideration may indeed create a more complex picture in terms of feature selection. The top 5 predictors in terms of selection frequency across forecasting samples are listed in Table 4. Regardless of the forecasting horizon $h$, the top predictor for ARMAr-LAS is the Goods Index. The other top predictors, also in the HCPI domain, include EA measurements (e.g., Services Index), or are specific to France and Germany (e.g., All-Items). In summary, ARMAr-LAS exploits cross-sectional information mainly focusing on prices, and accrues a forecasting advantage – as LAS uses many more variables to produce significantly worse forecasts.
Table 4: Five most frequently selected predictors. Selection percentages are ratios between the times a predictor appears in a forecast and the total number of forecasts (120 for $h$=12 and 96 for $h=24$).
Rank Selected Variables
$h$=12 $h=24$
Goods, Index Goods, Index
85.8% 85.4%
II° Industrial Goods, Index Services, Index
47.5% 43.8%
III° Services, Index All-Items (De)
40.8% 35.4%
32.5% 32.3%
All-Items (Fr) Industrial Goods, Index
24.2% 30.2%
@@ 5 Concluding Remarks
In this paper, we demonstrated that the probability of spurious correlations between stationary orthogonal or weakly correlated processes depends not only on the sample size, but also on the degree of predictors serial correlation. Through this result, we pointed out that serial correlation negatively affects the estimation and forecasting error bounds of LASSO. In order to improve the performance of LASSO in a time series context, we proposed an approach based on applying LASSO to pre-whitened (i.e., ARMA filtered) time series. This proposal relies on a working model that mitigates large spurious correlation and improves both estimation and forecasting accuracy. We characterized limiting distribution and feature selection consistency, as well as forecast and estimation error bounds, for our proposal. Furthermore, we assessed its performance through Monte Carlo simulations and an empirical application to Euro Area macroeconomic time series. Through simulations, we observed that ARMAr-LASSO, i.e., LASSO applied to ARMA residuals, reduces the probability of large spurious correlations and outperforms other LASSO-based methods from the literature in terms of both coefficient estimation and forecasting. The empirical application confirms that ARMAr-LASSO improves forecasting performance and produces more parsimonious models.
Based on the results obtained so far, we envision several avenues for future work. For instance, it would be of interest to derive the rate at which the distribution of the sample correlation coefficient approaches $\mathcal{D}(r)$, thus formalizing what we observed numerically (see Theorem 1 and Figure 1). From a more practical perspective, it would be of interest to develop guidelines on the number of response lags to include as additional regressors in the working model of ARMAr-LASSO. Also, other approaches could be explored for the tuning of $\lambda$ as alternatives to the BIC (see, e.g., the proposal by Adamek et al. (2023)). Finally, we intend to explore additional econometric applications; for instance, the analysis of EA macroeconomic data presented here could be replicated on other data sets, such as the FRED-MD dataset for the U.S.
Acknowledgements
Appendix A Upper Bound for $\psi_{min}$
To support this argument, we start by recalling an inequality that links off-diagonal elements and eigenvalues of $\widehat{\mathbf{C}}_{x}$; namely, $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ Because of this, for any given $\tau\in[0,1)$ we have
which emphasizes how the probability of a generic sample correlation being large in absolute value affects the probability of the minimum eigenvalue being small – and thus the estimation error bounds of the LASSO, as established by Bickel et al. (2009). As the next example shows, point the inequality $\widehat{\psi}_{min}^{x}\leq 1-\underset{i\neq j}{\text{max}}|\widehat{c}_{ij}^{x}|$ can be easily fixed.
Example A.1:
Let $e_{i}$ and $e_{j}$ be vectors from the standard basis of $\mathbb{R}^{n}$, $i,j\in{1,\dots,N}$. Moreover, let $x_{\pm}=2^{-1/2}(e_{i}\pm e_{j})$, satisfying $||x_{\pm}||_{2}=1$, and let $A$ be a correlation matrix with $a_{k}$ be the $k$-th column. Then we have
Thus, $\psi_{min}\leq 1-|a_{ij}|$ for all $i\neq j$ and so
$\psi_{min}\leq 1-\max_{i\neq j}|a_{ij}|\ \ .$
Appendix B Theoretical Results
B.1 Proof of Theorem 1
Consider a first order $n$-variate autoregressive process $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, where $\boldsymbol{\phi}$ is the $n\times n$ diagonal matrix with $\text{diag}(\boldsymbol{\phi})=(\phi_{1},\dots,\phi_{n})$ such that $|\phi_{i}|<1$ for each $i=1,\dots,n$, and the autoregressive residuals are assumed to be $\mathbf{u}_{t}\sim N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\boldsymbol{I}_{n}}\right)$. Here $\mathbf{x}_{0}\sim$ $N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}\right)$ and $\mathbf{x}_{t}\sim$ $N\mathopen{}\mathclose{{}\left(\boldsymbol{0}_{n},\mathbf{C}_{x}}\right)$ with $\mathopen{}\mathclose{{}\left(\mathbf{C}_{x}}\right)_{ii}=\frac{1}{1-\phi_{i}^{2}}$, and $\mathopen{}\mathclose{{}\left(\mathbf{C}_{x}}\right)_{ij}={c}_{ij}^{x}=0$, for $i\neq j$.
In this setting, we focus on the probability density of $\widehat{c}_{ij}^{x}$. We shall consider
$r=\frac{a_{ij}}{\sqrt{a_{ii}}\sqrt{a_{22}}}\ \ ,$
where $a_{i,j}=\sum_{t=1}^{T}(x_{i,t}-\overline{x}_{i})(x_{j,t}-\overline{x}_{j})$. In particular, when $c_{ij}^{u}=0$, $b=a_{ji}/a_{ii}$ and $v=a_{jj}-a_{ji}^{2}/a_{ii}$, then
$\cfrac{\sqrt{a_{ii}}\ b}{\sqrt{v}}=\cfrac{a_{ij}/\sqrt{a_{ii}a_{jj}}}{\sqrt{1-a_{ij}^{2}/(a_{ii}a_{jj})}}=\cfrac{r}{\sqrt{1-r^{2}}}\ \ .$ (B.7)
@@ B.1 Proof of Theorem 1
Note that $b$ is the least squares regression coefficient of $x_{jt}$ on $x_{it}$, and $v$ is the sum of the square of residuals of such regression. Thus, to derive the probability density of $r$ we need the probability densities of $b$ and $v$.
Probability Distribution of $b$.
@@
We start by deriving the probability distribution of $b$, the OLS regression coefficient for $x_{j}$ on $x_{i}$. The same holds if we regress $x_{i}$ on $x_{j}$.
Lemma B.1:
$b\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}\right)$.
Proof: We first focus on the distribution of the sample covariance between $x_{i}$ and $x_{j}$, which is
$\widehat{Cov}(x_{i},x_{j})=\frac{a_{ij}}{(T-1)}=\\ =\mathopen{}\mathclose{{}\left[\sum_{l=1}^{T-2}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\sum_{l=1}^{T-2}\phi_{j}^{l}\widehat{Cov}(u_{j[-l]},u_{i})+\widehat{Cov}(u_{i},u_{j})}\right](1-\phi_{i}\phi_{j})^{-1}\ \ ,$ (B.8)
where $\widehat{Cov}(u_{i[-l]},u_{j})=\sum_{t=l+1}^{T}(u_{i,t-l}-\overline{u}_{i})(u_{j,t}-\overline{u}_{j})/(T-l)$, for $i\neq j$, $\overline{u}_{i}=\frac{1}{T-l-1}\sum_{t=l+1}^{T}u_{i,t-l}$ and $\overline{u}_{j}=\frac{1}{T-l-1}\sum_{t=l+1}^{T}u_{j,t}$. Since $u_{i}$ and $u_{j}$ are standard Normal, we have (see  Glen et al., 2004 and Supplement K)
$\widehat{Cov}(u_{i},u_{j})\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0\>,\>\cfrac{1}{(T-1)}}\right)\ \ .$
Moreover, the quantity
$\eta_{ij}=\sum_{l=1}^{T-2}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\sum_{l=1}^{T-2}\phi_{j}^{l}\widehat{Cov}(u_{j[-l]},u_{i})\ \ ,$ (B.9)
is a linear combination of the sample covariances between the residual of a time series at time $t$ and the lagged residuals of the other time series. Note that $\eta_{ij}$ is a linear combination of $N\mathopen{}\mathclose{{}\left(0,\frac{\phi_{i}^{2l}}{T-l-1}}\right)$ and $N\mathopen{}\mathclose{{}\left(0,\frac{\phi_{j}^{2l}}{T-l-1}}\right)$. However, as $T\rightarrow\infty$, all the covariances in (B.9) converge to centred Normals with variance $\frac{1}{T-1}$. Therefore, because of $|\phi_{i}|,|\phi_{j}|<1$, for the convergence of the sum of infinite terms of geometric series, and by considering that $E\mathopen{}\mathclose{{}\left(\sum_{l=1}^{T-2}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})\sum_{l=1}^{T-2}\phi_{j}^{l}\widehat{Cov}(u_{j[-l]},u_{i})}\right)=0$, we have
@@
$Var\mathopen{}\mathclose{{}\left(\eta_{ij}}\right)=\frac{\phi_{i}^{2}}{(T-1)(1-\phi_{i}^{2})}+\frac{\phi_{j}^{2}}{(T-1)(1-\phi_{j}^{2})}=\frac{\phi_{i}^{2}+\phi_{j}^{2}-2\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})}\ \ .$
Therefore,
@@
$\eta_{ij}\stackrel{{\scriptstyle d}}{{\rightarrow}}N\mathopen{}\mathclose{{}\left(0,\frac{\phi_{i}^{2}+\phi_{j}^{2}-2\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})}}\right)$

### v3

@@ 4.2 Empirical Application
ARMAr-LAS FarSel 0.71*** 0.73*** 6.0||77.2 6.2||72.5
ARMAr-LAS AR(pp) 0.94 0.89*
LAS AR(pp) 1.36 1.08
GLS-LAS AR(pp) 1.43 1.44
ARDL-LAS AR(pp) 1.53 1.07
FarSel AR(pp) 1.32 1.21
Figure 4 summarizes patterns of selected predictors over time for LAS and ARMAr-LAS. The heatmaps represent the number of selected variables categorized according to the nine main domains (see above). LAS selects predictors largely, though not exclusively, from the domains ICS, M&IR and HUR. ARMAr-LAS is more targeted, selecting predictors almost exclusively in the HCPI domain (in Supplement F, we report the top 5 predictors in terms of selection frequency across forecasting samples). Note, however, that in a few instances (3 for $h=12$ and 2 for $h=24$) ARMAr-LAS does select many more predictors across multiple groups. Interestingly, these correspond to the period of the financial crisis (between the end of 2008 and the beginning of 2010), when negative shocks in some of the variables under consideration may indeed create a more complex picture in terms of feature selection. In summary, ARMAr-LAS exploits cross-sectional information mainly focusing on prices, and accrues a forecasting advantage – as LAS uses many more variables to produce significantly worse forecasts.
@@ 5 Concluding Remarks
In this paper, we demonstrated that the probability of spurious correlations between stationary orthogonal or weakly correlated processes depends not only on the sample size, but also on the degree of predictors serial correlation. Through this result, we pointed out that serial correlation negatively affects the estimation and forecasting error bounds of LASSO. In order to improve the performance of LASSO in a time series context, we proposed an approach based on applying LASSO to pre-whitened (i.e., ARMA filtered) time series. This proposal relies on a working model that mitigates large spurious correlation and improves both estimation and forecast accuracy. We characterized limiting distribution and feature selection consistency, as well as forecast and estimation error bounds, for our proposal. Furthermore, we assessed its performance through Monte Carlo simulations and an empirical application to Euro Area macroeconomic time series. Through simulations, we observed that ARMAr-LASSO, i.e., LASSO applied to ARMA residuals, reduces the probability of large spurious correlations and outperforms other LASSO-based methods from the literature in terms of both coefficient estimation and forecasting. The empirical application confirms that ARMAr-LASSO improves forecasting performance and produces more parsimonious models.
Based on the results obtained so far, we envision several avenues for future work. For instance, it would be of interest to derive the rate at which the distribution of the sample correlation coefficient approaches $\mathcal{D}(r)$, thus formalizing what we observed numerically in Figure 2. Another promising avenue for future work is the development of a desparsified ARMAr-LASSO to enable valid inference in high-dimensional time series with serially correlated predictors and error terms. Such an extension could build on the Bartlett-kernel Newey-West long run covariance estimator and could be compared with the recent inferential frameworks proposed by Chernozhukov2021 and Babii2024. We also note that the density in Proposition 1 provides a theoretical foundation for further advancement in testing correlations that link autoregressive processes.
Finally, we intend to explore additional econometric applications; for instance, the analysis of EA macroeconomic data presented here could be replicated on other data sets, such as the FRED-MD dataset for the U.S.
Appendix A Proofs
In this section, we provide the proof of our theoretical results.
A.1 Proof of Proposition 1
Let $\mathbf{x}_{t}=\boldsymbol{\phi}\mathbf{x}_{t-1}+\mathbf{u}_{t}$, $t=1,\dots,T$, be a first order $n$-variate autoregressive process as in Section 2.1. We focus on the probability density of $\widehat{c}_{ij}^{x}$. Following Anderson03 ch. 4, we shall consider $r=\frac{a_{ij}}{\sqrt{a_{ii}}\sqrt{a_{jj}}}$, where $a_{ij}=\sum_{t=1}^{T}(x_{i,t}-\overline{x}_{i})(x_{j,t}-\overline{x}_{j})$. In particular, when $c_{ij}^{u}=0$, $b=a_{ji}/a_{ii}$ and $v=a_{jj}-a_{ji}^{2}/a_{ii}$,
$\cfrac{\sqrt{a_{ii}}\ b}{\sqrt{v}}=\cfrac{a_{ij}/\sqrt{a_{ii}a_{jj}}}{\sqrt{1-a_{ij}^{2}/(a_{ii}a_{jj})}}=\cfrac{r}{\sqrt{1-r^{2}}}\ \ .$ (1)
@@ A.1 Proof of Proposition 1
Note that $b$ is the least squares regression coefficient of $x_{jt}$ on $x_{it}$, and $v$ is the sum of the square of residuals of such regression. Thus, according to (1), to obtain the probability density of $\widehat{c}_{ij}^{x}$, we need the distributions of $b$ and $v$.
Remark 1:
In contrast to asymptotic statements, our theoretical analysis is intended to derive distributions and densities of estimators that hold for finite $T$ as in Proposition 1. Hence, we will not employ the usual concepts of convergence in probability and in distribution; rather, we will use a notion of approximation, whose “precision” has been numerically evaluated in Section 2.1 ans Supplement B.
Throughout, the symbol $\overset{\mathrm{BE}}{\approx}$ denotes a finite–sample Gaussian approximation justified by Berry–Esseen type bounds and assessed via Monte Carlo simulations; it is not an asymptotic statement. The symbol $\approx$ is used exclusively to indicate a classical numerical approximation and does not refer to any asymptotic notion of convergence.
@@
Distribution of $b$. We start by deriving the sample distribution of $b$, the OLS regression coefficient for $x_{j}$ on $x_{i}$. The same holds if we regress $x_{i}$ on $x_{j}$.
Lemma 1:
@@
The sample probability distribution of $b$ is approximately $N\mathopen{}\mathclose{{\left(0,\frac{(1-\phi_{i}^{2}\phi_{j}^{2})(1-\phi_{i}^{2})}{(T-1)(1-\phi_{j}^{2})(1-\phi_{i}\phi_{j})^{2}}}}\right)$.
Proof of Lemma 1 We first focus on the distribution of the sample covariance between $x_{i,t}$ and $x_{j,t}$. Let $\widehat{Cov}(x_{i[-l]},x_{j})=\sum_{t=l+1}^{T}(x_{i,t-l}-\overline{x}_{i[-l]})(x_{j,t}-\overline{x}_{j})/(T-l-1)$, for $i\neq j$, where $\overline{x}_{i[-l]}=\frac{1}{T-l}\sum_{t=l+1}^{T}x_{i,t-l}$ and $\overline{x}_{j}=\frac{1}{T-l}\sum_{t=l+1}^{T}x_{j,t}$. By considering the AR(1) decomposition of $x_{i,t}$ and $x_{j,t}$, we have
$\frac{a_{ij}}{(T-1)}=\widehat{Cov}(x_{i},x_{j})\\ =\phi_{i}\phi_{j}\widehat{Cov}(x_{i[-1]},x_{j[-1]})+\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})+\phi_{j}\widehat{Cov}(x_{j[-1]},u_{i})+\widehat{Cov}(u_{i},u_{j})\ \ .$ (2)
Note that by moving $\phi_{i}\phi_{j}\widehat{Cov}(x_{i[-1]},x_{j[-1]})$ on the left side of the equality and adding and removing the quantity $\phi_{i}\phi_{j}\widehat{Cov}(x_{i},x_{j})$, after few algebra we obtain
$(1-\phi_{i}\phi_{j})\widehat{Cov}(x_{i},x_{j})=\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})+\phi_{j}\widehat{Cov}(x_{j,[-1]},u_{i})+\widehat{Cov}(u_{i},u_{j})\\ -\phi_{i}\phi_{j}\mathopen{}\mathclose{{\left(\widehat{Cov}(x_{i},x_{j})-\widehat{Cov}(x_{i[-1]},x_{j[-1]})}}\right)\ \ .$ (3)
Note that,
$\phi_{i}\widehat{Cov}(x_{i[-1]},u_{j})=\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\phi_{i}^{T-2}\widehat{Cov}(x_{i[-(T-2)]},u_{j}).$
The remainder terms $\phi_{i}^{T-2}\widehat{Cov}(x_{i[-(T-2)]},u_{j})$, $\phi_{j}^{T-2}\widehat{Cov}(x_{j[-(T-2)]},u_{i})$, and $\phi_{i}\phi_{j}\mathopen{}\mathclose{{\left(\widehat{Cov}(x_{i},x_{j})-\widehat{Cov}(x_{i[-1]},x_{j[-1]})}}\right)$ are negligible, and we may write
$\widehat{Cov}(x_{i},x_{j})\approx\\ \mathopen{}\mathclose{{\left[\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i[-l]},u_{j})+\sum_{l=1}^{T-3}\phi_{j}^{l}\widehat{Cov}(u_{j[-l]},u_{i})+\widehat{Cov}(u_{i},u_{j})}}\right](1-\phi_{i}\phi_{j})^{-1}\ \ .$ (4)
We can write $\widehat{Cov}(u_{i},u_{j})=\frac{1}{T-1}\sum_{t=1}^{T}u_{i,t}u_{j,t}-\frac{T}{T-1}\overline{u}_{i}\overline{u}_{j}$, where $\frac{T}{T-1}\overline{u}_{i}\overline{u}_{j}$ is negligible. Let $W_{t}\coloneq u_{i,t}u_{j,t}$, and $S_{T}\coloneq\frac{1}{\sqrt{T-1}}\sum_{t=1}^{T}W_{t}$. By the Berry–Esseen theorem berry1941; esseen1942, we have
$\sup_{a}\mathopen{}\mathclose{{\left|P\!\mathopen{}\mathclose{{\left(S_{T}\leq a}}\right)-\Phi(a)}}\right|\;\leq\;\frac{C\,E(|W_{1}|^{3})}{\sqrt{T-1}}\ \ ,$
for some universal constant $C<0.5$, and where $\Phi$ is the cumulative distribution function of the standard normal distribution. Moreover, $E(|W_{1}|^{3})=E(|u_{i,1}u_{j,1}|^{3})=E(|u_{i,1}|^{3})^{2}=E(|u_{j,1}|^{3})^{2}=(2\sqrt{2/\pi})^{2}=8/\pi\approx 2.546$. Thus, $\frac{C\,E(|W_{1}|^{3})}{\sqrt{T-1}}\;<\frac{0.5\times 2.55}{\sqrt{T-1}}\;=\;\frac{1.275}{\sqrt{T-1}}.$ Then, the Berry-Esseen theorem guarantees that
$\widehat{Cov}(u_{i,t},u_{j,t})\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0\>,\>\cfrac{1}{(T-1)}}}\right)\ \ .$
For a sample size $T>20$, the Berry-Esseen bound guarantees that $\frac{1.275}{\sqrt{T-1}}\;<\;0.292$. See Supplement G for numerical results.
Moreover, define
$\eta_{ij}=\sum_{l=1}^{T-3}\phi_{i}^{l}\widehat{Cov}(u_{i,t-l},u_{j,t})+\sum_{l=1}^{T-3}\phi_{j}^{l}\widehat{Cov}(u_{j,t-l},u_{i,t}).$ (5)
The quantity $\eta_{ij}$ is a linear combination of sample cross–covariances between the innovations of one series and lagged innovations of the other series. For each fixed lag $l$, by the Berry–Esseen theorem,
@@
$\widehat{Cov}(u_{i,t-l},u_{j,t})\overset{\mathrm{BE}}{\approx}N\!\mathopen{}\mathclose{{\left(0,\frac{1}{T-l-1}}}\right),$
and analogously for $\widehat{Cov}(u_{j,t-l},u_{i,t})$. Although the sample covariances at different lags are not independent, their mutual covariances are of smaller order with respect to the leading variance terms and are therefore negligible in the present approximation. Hence, exploiting $|\phi_{i}|,|\phi_{j}|<1$ and the convergence of the associated geometric series, we obtain
$\mathrm{Var}(\eta_{ij})\approx\sum_{l=1}^{T-3}\frac{\phi_{i}^{2l}}{T-l-1}+\sum_{l=1}^{T-3}\frac{\phi_{j}^{2l}}{T-l-1}\approx\frac{1}{T-1}\mathopen{}\mathclose{{\left(\frac{\phi_{i}^{2}}{1-\phi_{i}^{2}}+\frac{\phi_{j}^{2}}{1-\phi_{j}^{2}}}}\right).$
The approximation above treats the lag–specific sample cross–covariances as effectively uncorrelated when computing $\mathrm{Var}(\eta_{ij})$. Under temporally independent Gaussian innovations, the remaining cross–lag covariance terms arise only from finite–sample index overlap and contribute at a smaller order (typically $O(T^{-1})$) relative to the leading variance terms; we therefore neglect them as a second–order effect. Therefore,
$\eta_{ij}\overset{\mathrm{BE}}{\approx}N\mathopen{}\mathclose{{\left(0,\frac{\phi_{i}^{2}+\phi_{j}^{2}-2\phi_{i}^{2}\phi_{j}^{2}}{(T-1)(1-\phi_{i}^{2})(1-\phi_{j}^{2})}}}\right)\ \ ,$